# Log-concavity And The Multiplicative Properties of Restricted Partition Functions

### v1

@@ 5.1. The $m$-ary Partition Function
In 2019, Ganguly [18] showed that $b^{m}(n)$ is neither log-concave nor log-convex. Yet, for each $m$, the sequence $\{b^{m}(n)\}_{n\in\mathbb{N}}$ appears to be abundant with only finitely many instances of equality or failure.
@@
For each $m$, the sequence given by $\{b^{m}(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $b^{m}(\alpha)b^{m}(\beta)~{}\geq~{}b^{m}(\alpha+\beta)$ with instances of equality or failure as given by Table 4.
@@
$b^{m}(n)$ Equality at $(a,b)$ Failure at $(\alpha,\beta)$
$b^{2}(n)$ $(2,2),(3,2),(3,9)$ $a,b\leq 1$,
$(3,3),(5,3),(7,3)$
$b^{3}(n)$ $(7,8),(8,8)$ $a,b\leq 2$,
$(4,5),(4,8),(5,5),(5,7),(5,8)$
$b^{4}(n)$ $(5,7),(5,11),(5,15),(6,6)$ $a,b\leq 3$,
$(6,7),(6,10),(6,11),(6,14)$
$(6,15),(7,7),(7,9),(7,10)$
$(7,11),(7,13),(7,14),(7,15)$
And for $m\geq 5$, Failure only occurs for $\alpha,\beta\leq m$. For $m=5$, equality occurs for pairs $(\alpha,\beta)$ where $12m-2\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$. For $m\geq 6$, equality occurs for pairs $(\alpha,\beta)$ where $12m-3\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$.
For example,
$b^{5}(n)\ \text{has equality at}$ $b^{6}(n)\ \text{has equality at}$
$(8,7),(8,8)$ $(9,9)$
$(9,6),(9,7),(9,8),(9,9)$ $(10,8),(10,9),(10,10)$
$(11,7),(11,8),(11,9),(11,10),(11,11)$
@@ Acknowledgements
The authors would like to extend sincere gratitude to the anonymous referee, whose keen observations and criticism throughout the review process has greatly improved the quality of this paper.

### v4

@@ 5.1. The $m$-ary Partition Function
In 2019, Ganguly [19] showed that $b^{m}(\alpha)$ is neither log-concave nor log-convex. Yet, for each $m$, the sequence $\{b^{m}(\alpha)\}_{\alpha\in\mathbb{N}}$ appears to be abundant with instances of equality or failure according to congruence classes for $\alpha<m$ and only finitely many instances of equality (and no instances of failure) for $\alpha\geq m$. A regular pattern appears for $m\geq 6$.
@@
For each $2\leq m\leq 5$, the sequence given by $\{b^{m}(\alpha)\}_{\alpha\in\mathbb{N}}$ satisfies the inequality $b^{m}(\alpha)b^{m}(\beta)~{}\geq~{}b^{m}(\alpha+\beta)$ with instances of equality or failure as given by Table 4.
[…]
@@ Acknowledgements
The author would like to extend sincere gratitude to the anonymous referee, whose keen observations and criticism throughout the review process has greatly improved the quality of this paper.